Find the area of the region bounded by the curves

y=x^2-4x+4 and y=10-x^2

from x=10 and x=4


1
Expert's answer
2021-09-13T05:25:57-0400

Given curves are "y=x^2-4x+4 ; y=10-x^2"

Graph of the above curves are:






So, area bounded by these curves from "x=4" to "x=10" will be given as:

"∫_4^{10} (x^2-4x+4) dx +|∫_4^{10} (10-x^2)dx| \\
=∫_4^{10} (x-2)^2 dx + |[10x-\frac{x^3}{3}]_4^{10}| \\
= [\frac{(x-2)^3}{3}]_4^{10} + |100-\frac{1000}{3}-40+\frac{64}{3}| \\
=\frac{512}{3}-\frac{64}{3}+|60-\frac{936}{3}|\\
=\frac{448}{3}+\frac{756}{3}\\
=\frac{1204}{3}\\
=401.33 \ sq. units"